# Hyperbolic Network Latent Space Model with Learnable Curvature

### v1

@@ 6 Real-world Data Example
In this section, we employ the proposed hyperbolic latent space model to analyze the Simmons College Facebook friendship network (Traud et al.,, 2012). Within this anonymized network, individual nodes represent students, and edges connect nodes when the corresponding students are mutual friends on Facebook. The dataset offers supplementary details about each student, including their graduation year. To align with Ma et al., (2020), we preprocess the dataset, focusing solely on the largest connected component of students who graduated between 2006 and 2009. Subsequent to preprocessing, the network encompasses $1158$ nodes and $24449$ undirected edges.
For visualization purposes, we begin by fitting the $2$-dimensional latent space models. The estimated curvature parameter is $\hat{K}=1.450$, along with a log-likelihood value of $-149129.6$. The outcome of the testing procedure leads to the rejection of the null hypothesis that the underlying latent space geometry is Euclidean at a $5\%$ significance level. Moreover, the $95\%$ confidence interval for the latent space curvature is $(-1.532,-1.352)$. For comparison, we also fit a $2$-dimensional hyperbolic model with a standard curvature parameter of $K=1$, as well as a $2$-dimensional Euclidean model with a curvature of 00. The log-likelihoods of these two models are $-157616.8$ and $-166412.0$, respectively. This indicates two key observations: (1) hyperbolic latent space models provide a better fit for the network than the Euclidean model; and (2) selecting an appropriate hyperbolic space curvature can further enhance model fitting. Notably, in terms of parameter count, the $2$-dimensional hyperbolic models only introduce one additional curvature parameter when compared to their $2$-dimensional Euclidean counterpart.
@@
To further investigate the impact of the latent space dimension and curvature, we compare the network latent space models with different latent dimensions and curvatures with (1) classic model selection criteria BIC and AIC, calculated with the sample size of $\binom{1158}{2}$; and (2) average link prediction AUC scores, evaluated by randomly choosing $80\%$ entries of $A$ as the training set and evaluate AUC on the rest $20\%$ entries, repeated by $100$ times. The experiment results in Table 2 indicate that, within each dimension, the proposed hyperbolic model with a learnable curvature surpasses the performance of both the hyperbolic model with fixed $K=1$ and the Euclidean model. For estimating the dimension of the latent space, BIC prefers the $3$-dimensional model, while AIC is less conservative and favors the $5$-dimensional model. The highest average link prediction AUC score is achieved by the $6$-dimensional model.
@@
We also note the following observations: (1) Hyperbolic models are particularly powerful in low-dimensional settings. With an optimized curvature, a $3$-dimensional hyperbolic model can achieve similar link prediction performance of a $5$-dimensional Euclidean model. (2) The estimation of $K$ becomes smaller as the latent space dimension of the working model grows and the hyperbolic models have less improvement than the Euclidean model. Intuitively, this may be attributed to the trade-off between the latent space dimension and the curvature in achieving optimal model capacity. Increasing the dimension enhances the model’s capacity, while making the space more hyperbolic improves its ability to capture the network properties. Such intuition also explains why hyperbolic models with fixed $K=-1$ is worse than Euclidean models for $d\geq 5$. A similar finding can be found in Mathieu et al., (2019). (3) Nevertheless, for each $2\leq d\leq 7$, the proposed hyperbolic model with learnable curvature still outperforms the other two models for higher-dimensional cases for all four criteria reported in Table 2. And for all latent space dimensions, the bootstrap tests reject the null hypothesis of $K=0$. These observations shed light on the significance of accounting for the latent space curvature in network analysis.
Dimension Model Log-Likelihood (×$\times 10^{5}$) BIC (×$\times 10^{5}$) AIC (×$\times 10^{5}$) AUC (sd)
d = 2 $\mathbb{E}_{2}$1.664 3.639 3.375 0.855 (0.004)
$\mathbb{H}_{2}^{-1}$1.577 3.463 3.199 0.870 (0.002)
$\mathbb{H}_{2}^{-1.450}$1.491 3.293 3.029 0.884 (0.002)
d = 3 $\mathbb{E}_{3}$1.425 3.315 2.919 0.883 (0.002)
$\mathbb{H}_{3}^{-1}$1.358 3.183 2.787 0.903 (0.002)
$\mathbb{H}_{3}^{-1.283}$1.356 3.179 2.782 0.902 (0.002)
d = 4 $\mathbb{E}_{4}$1.343 3.308 2.779 0.895 (0.002)
$\mathbb{H}_{4}^{-1}$1.318 3.257 2.728 0.908 (0.002)
$\mathbb{H}_{4}^{-0.516}$1.304 3.230 2.702 0.909 (0.002)
d = 5 $\mathbb{E}_{5}$1.306 3.389 2.728 0.903 (0.002)
$\mathbb{H}_{5}^{-1}$1.310 3.397 2.736 0.910 (0.002)
$\mathbb{H}_{5}^{-0.272}$1.285 3.349 2.687 0.911 (0.002)
d = 6 $\mathbb{E}_{6}$1.291 3.513 2.720 0.907 (0.002)
$\mathbb{H}_{6}^{-1}$1.306 3.544 2.751 0.911 (0.002)
$\mathbb{H}_{6}^{-0.170}$1.275 3.483 2.689 0.912 (0.002)
d = 7 $\mathbb{E}_{7}$1.279 3.645 2.720 0.909 (0.002)
$\mathbb{H}_{7}^{-1}$1.303 3.694 2.768 0.911 (0.002)
$\mathbb{H}_{7}^{-0.121}$1.269 3.625 2.699 0.912 (0.002)
@@
In conclusion, our data analysis showcases the significant advantages of the proposed hyperbolic latent space model. It not only offers better model fitting and visualization capabilities but also enhances downstream task performance, demonstrating its potential as a valuable tool for analyzing and understanding complex network data.
@@ 7 Discussion
Our work raises several promising directions for future research. From a modeling perspective, one may include additional node and edge parameters in Equation 4 to further incorporate node-level or edge-wise information. For example, similar to Ma et al., (2020) we can let $P_{ij}=\sigma(d_{\mathbb{H}_{d}^{-K}}(z_{i},z_{j})+\alpha_{i}+\alpha_{j}+\beta x_{ij})$, with $\alpha_{i},1\leq i\leq n$ accounting for node heterogeneity and $x_{ij},1\leq i<j\leq n$ representing edge-related covariates. The node parameters $\alpha_{i}$ could be unbounded to better model the sparse network data. These extensions that incorporate more realistic settings are worthwhile for future exploration. Another interesting generalization of the proposed model is to consider latent spaces with non-constant local curvatures. Taking social networks as an example, the local curvature near the latent positions of hub nodes might be larger such that the pairwise distances on the latent space are shorter among these nodes. It is also interesting to further investigate the model selection problem on the interplay between latent space curvature and the dimension, which relates to a technical problem of generalizing Theorem 4.1 towards higher-dimensional hyperbolic spaces.
Beyond network data, hyperbolic latent space models or embedding methods can be used for modeling tree structured data (e.g., phylogenetic tree; Sala et al.,, 2018) and other heterogeneous relational data (e.g., text data parsed as graphs; Tifrea et al.,, 2018). With the insights provided by Theorem 4.1, it would be interesting to investigate whether including the learnable curvature parameter can help increase the downstream task performance of these general hyperbolic embedding methods.

### v2

@@ 6 Real-world Data Example
In this section, we employ the proposed hyperbolic latent space model to analyze the Simmons College Facebook friendship network (Traud et al., 2012). Within this anonymized network, individual nodes represent students, and edges connect nodes when the corresponding students are mutual friends on Facebook. The dataset offers supplementary details about each student, including their graduation year. To align with Ma et al. (2020), we preprocess the dataset, focusing solely on the largest connected component of students who graduated between 2006 and 2009. Subsequent to preprocessing, the network encompasses $1158$ nodes and $24449$ undirected edges.
For visualization purposes, we begin by fitting the $2$-dimensional latent space models. The estimated curvature parameter is $\hat{K}=1.450$, along with a log-likelihood value of 149129.6-149129.6. The outcome of the testing procedure leads to the rejection of the null hypothesis that the underlying latent space geometry is Euclidean at a $5\%$ significance level. Moreover, the $95\%$ confidence interval for the latent space curvature is (1.532,1.352)(-1.532,-1.352). For comparison, we also fit a $2$-dimensional hyperbolic model with a standard curvature parameter of $K=1$, as well as a $2$-dimensional Euclidean model with a curvature of 0. The log-likelihoods of these two models are 157616.8-157616.8 and 166412.0-166412.0, respectively. This indicates two key observations: (1) hyperbolic latent space models provide a better fit for the network than the Euclidean model; and (2) selecting an appropriate hyperbolic space curvature can further enhance model fitting. Notably, in terms of parameter count, the $2$-dimensional hyperbolic models only introduce one additional curvature parameter when compared to their $2$-dimensional Euclidean counterpart.
@@
To further investigate the impact of the latent space dimension and curvature, we compare network latent space models with different latent dimensions and curvatures with (1) classic model selection criteria BIC and AIC, calculated with the sample size of $\binom{1158}{2}$; and (2) average link prediction AUC scores, evaluated by randomly choosing $80\%$ entries of $A$ as training set and evaluate AUC in the rest of $20\%$ entries, repeated $100$ times. The results of the experiment in Table 1 indicate that, within each dimension, the proposed hyperbolic model with a learnable curvature exceeds the performance of both the hyperbolic model with fixed $K=1$ and the Euclidean model. For estimating the dimension of the latent space, BIC prefers the $3$-dimensional model, while AIC is less conservative and favors the $5$-dimensional model. The highest average link prediction AUC score is achieved by the $6$-dimensional model. In the Supplementary Materials, we further compare with the link prediction performance of the Poincaré-embedding method proposed in Nickel and Kiela (2017) and demonstrate the advantages of the proposed model, particularly in fitting the optimal curvature and employing robust initialization.
@@
We also note the following observations: (1) Hyperbolic models are particularly powerful in low-dimensional settings. With an optimized curvature, a $3$-dimensional hyperbolic model can achieve similar link prediction performance of a $5$-dimensional Euclidean model. (2) The estimation of $K$ becomes smaller as the latent space dimension of the working model grows and the hyperbolic models have less improvement than the Euclidean model. Intuitively, this may be attributed to the trade-off between the latent space dimension and the curvature in achieving optimal model capacity. Increasing the dimension enhances the model’s capacity, while making the space more hyperbolic improves its ability to capture the network properties. Such intuition also explains why hyperbolic models with fixed K=1K=-1 is worse than Euclidean models for $d\geq 5$. A similar phenomenon can be found in Mathieu et al. (2019). (3) Nevertheless, for each $2\leq d\leq 7$, the proposed hyperbolic model with learnable curvature still outperforms the other two models for higher-dimensional cases for all four criteria reported in Table 1. And for all latent space dimensions, the bootstrap tests reject the null hypothesis of $K=0$. These observations shed light on the significance of accounting for the latent space curvature in network analysis.
Dimension Model Log-Likelihood ($\times 10^{5}$) BIC ($\times 10^{5}$) AIC ($\times 10^{5}$) AUC (sd)
d = 2 $\mathbb{E}_{2}$ -1.664 3.639 3.375 0.855 (0.004)
$\mathbb{H}_{2}^{-1}$ -1.577 3.463 3.199 0.870 (0.002)
$\mathbb{H}_{2}^{-1.450}$ -1.491 3.293 3.029 0.884 (0.002)
d = 3 $\mathbb{E}_{3}$ -1.425 3.315 2.919 0.883 (0.002)
$\mathbb{H}_{3}^{-1}$ -1.358 3.183 2.787 0.903 (0.002)
$\mathbb{H}_{3}^{-1.283}$ -1.356 3.179 2.782 0.902 (0.002)
d = 4 $\mathbb{E}_{4}$ -1.343 3.308 2.779 0.895 (0.002)
$\mathbb{H}_{4}^{-1}$ -1.318 3.257 2.728 0.908 (0.002)
$\mathbb{H}_{4}^{-0.516}$ -1.304 3.230 2.702 0.909 (0.002)
d = 5 $\mathbb{E}_{5}$ -1.306 3.389 2.728 0.903 (0.002)
$\mathbb{H}_{5}^{-1}$ -1.310 3.397 2.736 0.910 (0.002)
$\mathbb{H}_{5}^{-0.272}$ -1.285 3.349 2.687 0.911 (0.002)
d = 6 $\mathbb{E}_{6}$ -1.291 3.513 2.720 0.907 (0.002)
$\mathbb{H}_{6}^{-1}$ -1.306 3.544 2.751 0.911 (0.002)
$\mathbb{H}_{6}^{-0.170}$ -1.275 3.483 2.689 0.912 (0.002)
d = 7 $\mathbb{E}_{7}$ -1.279 3.645 2.720 0.909 (0.002)
$\mathbb{H}_{7}^{-1}$ -1.303 3.694 2.768 0.911 (0.002)
$\mathbb{H}_{7}^{-0.121}$ -1.269 3.625 2.699 0.912 (0.002)
@@
In conclusion, our data analysis shows the significant advantages of the proposed hyperbolic latent space model. It not only offers better model fitting and visualization capabilities but also enhances downstream task performance, demonstrating its potential as a valuable tool for analyzing and understanding complex network data.
@@ 7 Discussion
Our work also raises several promising directions for future research. From a modeling perspective, one may include additional node and edge parameters in Equation 4 to further incorporate node-level or edge-wise information. For example, similar to Ma et al. (2020) we can let $P_{ij}=\sigma(d_{\mathbb{H}_{d}^{-K}}(z_{i},z_{j})+\alpha_{i}+\alpha_{j}+\beta x_{ij})$, with $\alpha_{i},1\leq i\leq n$ accounting for node heterogeneity and $x_{ij},1\leq i<j\leq n$ representing edge-related covariates. The node parameters $\alpha_{i}$ could be unbounded to better model the sparse network data. These extensions that incorporate more realistic settings are worthwhile for future exploration. Another interesting generalization of the proposed model is to consider latent spaces with non-constant local curvatures. Taking social networks as an example, the local curvature near the latent positions of hub nodes might be larger such that the pairwise distances on the latent space are shorter among these nodes. It is also interesting to further investigate the model selection problem on the interplay between latent space curvature and the dimension, which relates to a technical problem of generalizing Theorem 4.1 towards higher-dimensional hyperbolic spaces.
Beyond network data, hyperbolic latent space models or embedding methods can be used for modeling tree structured data (e.g., phylogenetic tree; Sala et al., 2018) and other heterogeneous relational data (e.g., text data parsed as graphs; Tifrea et al., 2018). With the insights provided by Theorem 4.1, it would be interesting to investigate whether including the learnable curvature parameter can help increase the downstream task performance of these general hyperbolic embedding methods.